# Do I Know This Entity? Knowledge Awareness and Hallucinations in Language Models

### v1

@@ 8 Related Work
Recent advances in mechanistic interpretability in language models (Ferrando et al., 2024) have shed light on the factual recall process in these systems. Key discoveries include the aggregation of entity tokens (Nanda et al., 2023), the importance of early MLPs for entity processing (Meng et al., 2022b), and the identification of specialized extraction relation heads (Geva et al., 2023; Chughtai et al., 2024). Despite these insights, there remains a significant gap in our understanding of the mechanisms underlying failures in attribute extraction leading to hallucinations. Gottesman & Geva (2024) demonstrated that the performance of probes trained on the residual streams of entities correlates with its ability to answer questions about them accurately. Yuksekgonul et al. (2024) established a link between increased attention to entity tokens and improved factual accuracy. (Yu et al., 2024) proposed two mechanisms for non-factual hallucinations: inadequate entity enrichment in early MLPs and failure to extract correct attributes in upper layers. Our research aligns with studies on hallucination prediction (Kossen et al., 2024; Varshney et al., 2023), particularly those engaging with model internals (CH-Wang et al., 2023; Azaria & Mitchell, 2023). Additionally, our work contributes to the growing body of literature on practical applications of sparse autoencoders, as investigated by Marks et al. (2024); Krzyzanowski et al. (2024). While the practical applications of sparse autoencoders in language model interpretation are still in their early stages, our research demonstrates their potential.
@@ 9 Conclusions
In this paper, we use sparse autoencoders to identify directions in the model’s representation space that encode a form of self-knowledge about entities. These directions, found in the base model, are causally relevant to the knowledge refusal behavior in the chat-based model. We demonstrated that, by manipulating these directions, we can control the model’s tendency to refuse answers or hallucinate information. We also provide insights into how the entity recognition directions influence the model behavior, such as regulating the attention paid to entity tokens, and their influence in expressing knowledge uncertainty. Finally, we uncover directions representing model uncertainty to specific queries, capable of discriminating between correct and mistaken answers. This work contributes to our understanding of language model behavior and opens avenues for improving model reliability and mitigating hallucinations.
@@ Acknowledgments
This work was conducted as part of the ML Alignment & Theory Scholars (MATS) Program. We want to express our sincere gratitude to McKenna Fitzgerald for her guidance and support during the program, to Matthew Wearden for his thoughtful feedback on the manuscript, and to Wes Gurnee for initial discussions that helped shape this work. We want to extend our gratitude to Adam Karvonen, Can Rager, Bart Bussmann, Patrick Leask and Stepan Shabalin for the valuable input during MATS. Lastly, we thank the entire MATS and Lighthaven staff for creating the environment that made this research possible. Javier Ferrando is supported by the Spanish Ministerio de Ciencia e Innovación through the project PID2019-107579RB-I00 / AEI / 10.13039/501100011033. Portions of this work were supported by the Long Term Future Fund.
@@ Appendix I Activation Patching on Gemma 2 2B
Appendix M Change in Attention Scores to Entities After Steering
Gemma 2 2B (Figures 16 and 17) and Gemma 2 9B (Figures 18 and 19) average attention scores to entity tokens after steering with the top known entity latents and top unknown entity latents. Error bars indicate standard deviation. For the known entity latent steering we use prompts with unknown entities, for the unknown entity latent steering we use prompts with known entities. The strength of the steering coefficient is . We show heads starting from layer 15 in Gemma 2 2B and layer 25 in Gemma 2 9B, coinciding with the point where information propagates to the last position.

### v2

@@ 8 Related Work
Recent advances in mechanistic interpretability in language models (Ferrando et al., 2024) have shed light on the factual recall process in these systems. Key discoveries include the aggregation of entity tokens (Nanda et al., 2023), the importance of early MLPs for entity processing (Meng et al., 2022b), and the identification of specialized extraction relation heads (Geva et al., 2023; Chughtai et al., 2024). Despite these insights, there remains a significant gap in our understanding of the mechanisms underlying failures in attribute extraction leading to hallucinations. Gottesman & Geva (2024) demonstrated that the performance of probes trained on the residual streams of entities correlates with the model’s ability to answer questions about them accurately. Yuksekgonul et al. (2024) established a link between increased attention to entity tokens and improved factual accuracy. (Yu et al., 2024) proposed two mechanisms for non-factual hallucinations: inadequate entity enrichment in early MLPs and failure to extract correct attributes in upper layers. Our research aligns with studies on hallucination prediction (Kossen et al., 2024; Varshney et al., 2023), particularly those engaging with model internals (CH-Wang et al., 2024; Azaria & Mitchell, 2023). Previous work has trained probes to predict truthfulness of the produced outputs (Li et al., 2023) with Joshi et al. (2024) showing this can be detected in activation space before the model generation, which can be related to our results on ‘uncertainty directions’ discovered in Section 7. Additionally, our work contributes to the growing body of literature on practical applications of sparse autoencoders, as investigated by Marks et al. (2024); Krzyzanowski et al. (2024). While the practical applications of sparse autoencoders in language model interpretation are still in their early stages, our research demonstrates their potential.
@@ 9 Conclusions
In this paper, we use sparse autoencoders to identify directions in the model’s representation space that suggest the presence of encoded knowledge awareness about entities. These directions, found in the base model, are causally relevant to the knowledge refusal behavior in the chat-based model. We demonstrated that, by manipulating these directions, we can control the model’s tendency to refuse answers or hallucinate information. We also provide insights into how the entity recognition directions influence the model behavior, such as regulating the attention paid to entity tokens, and their influence in expressing knowledge uncertainty. Finally, we uncover directions representing model uncertainty to specific queries, capable of discriminating between correct and mistaken answers. While our primary focus in this work centers on the representation of knowledge awareness and uncertainty, the methodology we present for discovering these latent directions is generalizable to any other type of binary (Section 3) and continuous (Section 7) features. This work contributes to our understanding of language model behavior and opens avenues for improving model reliability and mitigating hallucinations.
@@ Acknowledgments
This work was conducted as part of the ML Alignment & Theory Scholars (MATS) Program. We want to express our sincere gratitude to McKenna Fitzgerald for her guidance and support during the program, to Matthew Wearden for his thoughtful feedback on the manuscript, and to Wes Gurnee for initial discussions that helped shape this work. We want to extend our gratitude to Adam Karvonen, Can Rager, Bart Bussmann, Patrick Leask and Stepan Shabalin for the valuable input during MATS. Lastly, we thank the entire MATS and Lighthaven staff for creating the environment that made this research possible. Portions of this work were supported by the Long Term Future Fund. Javier Ferrando is supported by the fellowship within the “Generación D” initiative, Red.es, Ministerio para la Transformación Digital y de la Función Pública, for talent atraction (C005/24-ED CV1). Funded by the European Union NextGenerationEU funds, through PRTR. Lastly, we appreciate the anonymous reviewers for their useful comments.
@@ Appendix I Activation Patching on Gemma 2 2B
Appendix M Change of Attention Scores to Entities After Steering
Gemma 2 2B (Figures 16 and 17), Gemma 2 9B (Figures 18 and 19) and Llama 3.1 8B (Figures 20 and 21) average attention scores to entity tokens after steering with the top known entity latents and top unknown entity latents. Error bars indicate standard deviation. For the known entity latent steering we use prompts with unknown entities, for the unknown entity latent steering we use prompts with known entities. The strength of the steering coefficient is for Gemma models and for Llama 3.1 8B. We show heads starting from layer the latent is found, and the steering is done.
We have assessed the statistical significance of attention score changes by comparing steering with entity recognition latents versus with random SAE latents using the same layer and steering coefficient. We conduct t-tests where the null hypothesis states that both steerings would yield identical mean attention scores differences across downstream attention heads. The alternative hypothesis is that known entity latents would increase mean attention scores, while unknown entity latents would decrease them. We tested against 10 different random SAE latents using 100 distinct prompts.
Results indicate that for Gemma 2 2B, Gemma 2 9B and Llama 3.1 8B, steering with the top known entity latent shows statistically significant larger average attention score when compared to random SAE latents on 10/10, 10/10 and 7/10 cases respectively. When steering with the top unknown entity latent it shows statistically significant lower average attention score when compared to random SAE latents on 9/10, 1/10 and 10/10 cases respectively. As shown in Figure 19 (top), Gemma 2 9B top unknown entity latent doesn’t show strong reductions. However the second unknown entity latent shows significant differences in 9/10 tests.
@@
Appendix Q Llama 3.1 8B Replication
We extend our experimental analysis to Llama 3.1 8B (Grattafiori et al., 2024), using the SAEs suite from LlamaScope (He et al., 2024), which offers per-layer pretrained SAEs. Following the methodology described in Section 3, we detect both known and unknown entity latents within the model. The distribution of the Top 5 latents across layers (Figure 25) exhibit consistent patterns with previous findings, with the most effective and generalizable latent representations concentrated in the intermediate layers.
Steering experiments using the top unknown entity latent reveal increase refusal rates in the instruction-tuned model (Figure 26). Conversely, when we orthogonalize the model weights with respect to this direction, refusal rates decrease. Since the original model’s refusal rate on unknown entity prompts is high (Figure 26 left), we include the refusal rates on prompts with known entities (Figure 26 right). Notably, steering with the top known entity latent did not produce a corresponding decrease in refusals.
Further analysis reveal similar findings to those in Gemma regarding attention patterns: steering with the top known entity latent increases the attention scores to the entity (Figure 27 top), while unknown entity latent steering result in diminished attention scores (Figure 27 bottom).
The replication of our key findings—originally observed in Gemma—across Llama 3.1 8B strengthens our confidence in both our methodological approach and the broader applicability of our results. This generalization is particularly noteworthy given the substantial architectural differences between the two models and their respective SAEs.
Appendix R Activation Frequency on Songs Data After knowledge Cutoff
| Model | Known Entity Latent | Unknown Entity Latent |
| Gemma 2 2B | 6% | 53% |
| Gemma 2 9B | 22% | 55% |
| Llama 3.1 8B | 13.4% | 76% |
Appendix S Token Likelihood Hypothesis
An alternative explanation for our observed entity recognition latents is the token likelihood hypothesis: the latents might simply encode token likelihood rather than actual knowledge about entities. Under this hypothesis, when processing a token sequence (), the activation of our discovered latents at position could be explained by the model’s ability to predict token from previous context. For instance, given a well-known movie title, the model would more easily predict subsequent tokens, potentially triggering what we interpret as ‘known entity’ latents. Conversely, for unknown entities, lower token likelihood might activate our ‘unknown entity’ latents. This represents a plausible confounding factor, as tokens comprising well-known entity names are inherently more predictable in the training distribution than those of unknown entities.
To test this hypothesis, we analyze token likelihood on a broad text corpus from the FineWeb dataset (Penedo et al., 2024). For each token position , we compute both the entity recognition latent activations and the probability of the ground-truth token being predicted from position , . If the token likelihood hypothesis were true, we would expect strong correlations between these measures.
Our analysis reveals several key findings that challenge this hypothesis:
Entity recognition latents activate selectively, firing on only a small fraction of tokens (e.g. 0.6% for known entity latents and 0.5% for unknown entity latents in Gemma 2 2B, see Table 9).
The correlations between latent activations and token prediction probabilities are negligible across all tested models (Figure 28).
We explored various other potential relationships, including dependencies on the perplexity of surrounding tokens and next-token prediction entropy, finding no substantial correlations.
While we observe that tokens where unknown entity latents activate tend to have lower prediction probabilities compared to the baseline, this effect is modest given the latents’ sparse activation patterns. These findings suggest that token predictability alone cannot explain the behavior of our entity recognition latents, supporting our interpretation that they encode a more sophisticated form of knowledge awareness.
| Model | Latent | Activation Frequency | Correlation with |
| Gemma 2 2B | Known | 0.006 | 0.067 (p=0.000e+00) |
| Unknown | 0.005 | -0.000 (p=9.960e-01) | |
| Gemma 2 9B | Known | 0.009 | 0.062 (p=0.0e+00) |
| Unknown | 0.009 | 0.009 (p=1.045e-12) | |
| Llama 3.1 8B | Known | 0.002 | 0.003 (p=1.380e-02) |
| Unknown | 0.026 | 0.047 (p=0.000e+00) |